OpenStudy (anonymous):

create 2 problems two simplify multiplication and division with algebraic expressions

OpenStudy (anonymous):

ok lets do it!

OpenStudy (anonymous):

i am not exactly sure what "simplify multiplication and divsion" means you can multiply if you like, and you can divide or you can combine them

OpenStudy (anonymous):

for example \[(x^2+x+1)(x+2)=x^3+3 x^2+3 x+2\] is a multiplication

OpenStudy (anonymous):

oh like \[\frac{x+1}{x^2-4}\times \frac{x+2}{x^2-1}\] where you factor and cancel?

OpenStudy (anonymous):

yeah exactly as long as you can simplify it i think were good

OpenStudy (anonymous):

ok than we can use that example, because the denominator in both of those expressions is the difference of two square so you can write \[\frac{x+1}{x^2-4}\times \frac{x+2}{x^2-1}=\frac{(x+1)(x+2)}{(x+2)(x-2)(x+1)(x-1)}\] then cancel away

OpenStudy (anonymous):

\[\frac{\cancel{(x+1)}\cancel{(x+2)}}{\cancel{(x+2)}(x-2)\cancel{(x+1)}+(x-1)}=\frac{1}{(x-2)(x-1)}\]

OpenStudy (anonymous):

ooops that plus sign is a typo

OpenStudy (anonymous):

okay thats fine so which one is multpication, subtraction, division and addiction can you number them lol even though i know

OpenStudy (anonymous):

just four problems in all sub, and addit have to have a common denoimater

OpenStudy (anonymous):

that one above was both a multiplication and a division

OpenStudy (anonymous):

you want another just division?

OpenStudy (anonymous):

yes sure that would be great another one

OpenStudy (anonymous):

ok lets do just a division one since \((x^2+x+1)(x+2)=x^3+3 x^2+3 x+2\) if you write \[\frac{x^3+3x^2+3x+2}{x+2}\] you know the answer is \(x^2+x+1\)

OpenStudy (anonymous):

so a division only problem could be \[\frac{x^3+3x^2+3x+2}{x+2}=x^2+x+1\]

OpenStudy (anonymous):

now the subtraction and addition

OpenStudy (anonymous):

right you work it out by doing the multiplication in advance

OpenStudy (anonymous):

addition, can it be easy?

OpenStudy (anonymous):

ohhhhh okay thats even better

OpenStudy (anonymous):

\[\frac{x}{x+2}+\frac{3}{x+2}=\frac{x+3}{x+2}\] like that? or do the denominators have to be different?

OpenStudy (anonymous):

yes it can be super easy as long as it has a common deniomater

OpenStudy (anonymous):

no i like that one its perfect ( :

OpenStudy (anonymous):

ok that would be an addition one you can do a subtraction one just as easily put a minus sign instead of a plus sign

OpenStudy (anonymous):

or you could change it a bit, \[\frac{x}{x-3}-\frac{5}{x-3}=\frac{x-5}{x-3}\] or whatever

OpenStudy (anonymous):

just post the questions here and i will help

OpenStudy (anonymous):

and dont say it is a test, people get all upset

OpenStudy (anonymous):

okay im about to post it now sorry let me delete that then

OpenStudy (anonymous):

you can post right here if you like

OpenStudy (anonymous):

i will i just dont know how its like this graph hmmmmmm

OpenStudy (anonymous):

post a screen shot

OpenStudy (anonymous):

im trying to do that now lol

OpenStudy (anonymous):

can you see that im going to post the 3 question also

OpenStudy (anonymous):

1. find the perimeter of the mansion Remember to simplify your answer by combining like terms 2.How much larger is the perimeter of the Outdoor Pool than the perimeter of the Indoor Pool? 3.Find the volume of the Outdoor Pool.

OpenStudy (anonymous):

will be so helpful if you could do that( : im going to delte the other questions people be lurking

OpenStudy (anonymous):

ok let me see if i can enlarge it to read it

OpenStudy (anonymous):

okay ( :

OpenStudy (anonymous):

is that whole weird shape the mansion?

OpenStudy (anonymous):

yes lmao that is the mansion

OpenStudy (anonymous):

ok we can do this but it is going to take like forever we have to multiply and add a bunch of stuff

OpenStudy (anonymous):

okay well as long as you have the time im here so what stuff do we have to do explain slowly im not good with math weakest subject

OpenStudy (anonymous):

damn i got logged out sorry still there?

OpenStudy (anonymous):

its okay your here helping thats all that matters ( :

OpenStudy (anonymous):

lol ok you have two sides that are \(6x^2\) for a total of \(12x^2\) write that down somewhere we have a lot to go

OpenStudy (anonymous):

writing it now lol 6x2 and 12x2

OpenStudy (anonymous):

one side is \(3x^4-2\) and there are four of those for a total of \(12x^4-8\)

OpenStudy (anonymous):

just write down the \(12x^2\)

OpenStudy (anonymous):

wrote it down now

OpenStudy (anonymous):

one sides is \(4x^4+3x^3+x^2+4x\) and there are two of those for a total of \[8x^4+6x^3+2x^2+4x\] write that one down

OpenStudy (anonymous):

wrote it down now

OpenStudy (anonymous):

one side is \(3x^5+5x^4+x^3+2x^2+x+6\) and now we have to figure out how many of those there are

OpenStudy (anonymous):

if we ever get disconnected email me at doogil@ymail.com lol just incase

OpenStudy (anonymous):

so do i write that down to

OpenStudy (anonymous):

ok i am gong to say there are 4 of those for a total of \[12x^5+20x^4+4x^3+8x^2+4x+24\] write that one down

OpenStudy (anonymous):

just wrote it am i combing it with the others or just writing it separately?

OpenStudy (anonymous):

okay gothca i just wrote it down

OpenStudy (anonymous):

typo there \[8x^4+16x^4+16x^3+4x^2+20x\]

OpenStudy (anonymous):

one more to go one corner is \(4x\) and there are 8 of those for a total of \(32x\) now we can add

OpenStudy (anonymous):

hey you still here

OpenStudy (anonymous):

if i didn't make a mistake it will be a miracle this question really sucks now lets see what we get when we add

OpenStudy (anonymous):

okay so do i write that down ?

OpenStudy (anonymous):

yeah lets write it all down this is what i got

OpenStudy (anonymous):

okay so thats good then lol mistakes happen now lets add them up like you said ugggh

OpenStudy (anonymous):

\[12x^2+12x^4-8+8x^4+6x^3+2x^2+4x+8x^5+16x^4+16x^3+4x^2+20x+12x^5+20x^4+4x^3+8x^2+4x+24+32x\]

OpenStudy (anonymous):

oops too big to fit on one line no matter, we will add by machine

OpenStudy (anonymous):

lmaoooo okay yeah i can barely see the problem

OpenStudy (anonymous):

\[12x^2+12x^4-8+8x^4+6x^3+2x^2+4x\]\[+8x^5+16x^4+16x^3+4x^2+20x+12x^5+20x^4+4x^3+8x^2+4x+24+32x\]

OpenStudy (anonymous):

\[20 x^5+56 x^4+26 x^3+26 x^2+60 x+16\]

OpenStudy (anonymous):

thats exactly what i got for the first answer

OpenStudy (anonymous):

you gotta be kidding really?

OpenStudy (anonymous):

ok now i am exhausted what is the next question?

OpenStudy (anonymous):

lmaooooo yes im seriously i got it

OpenStudy (anonymous):

2.How much larger is the perimeter of the Outdoor Pool than the perimeter of the Indoor Pool? 3. Find the volume of the Outdoor Pool.

OpenStudy (anonymous):

jeez site is messing up on me keeps closing hold one

OpenStudy (anonymous):

omggg i didnt know you was sleeply okay lets hurry

OpenStudy (anonymous):

not sleepy, exhausted from this problem

OpenStudy (anonymous):

alright lets hurry upppp then sorry for the long problem lol

OpenStudy (anonymous):

that's ok i can email if i get stuck

OpenStudy (anonymous):

perimeter of outside pool is \[2(3x^5+2x)+2(3x^4-2x^2)\] lets cheat again

OpenStudy (anonymous):

lolllll yes lets cheat that sounds lovely ( :

OpenStudy (anonymous):

\[6 x^5+6 x^4-4 x^2+4 x\] http://www.wolframalpha.com/input/?i=2%283x^5%2B2x%29%2B2%283x^4-2x^2%29

OpenStudy (anonymous):

alright so thats the answer for the second, one ( :lets cheat for the 3rd two

OpenStudy (anonymous):

perimeter of inside pool is \[2(x^4+x^3-x)+2(2x^5+3x)\] oh no not yet it isn't!!

OpenStudy (anonymous):

ohhh damnnnn lollll well lets keep going

OpenStudy (anonymous):

we have to compute the perimeter of the inside pool, then subtract

OpenStudy (anonymous):

you didn't submit that answer did you?

OpenStudy (anonymous):

okay so our we done with the outside pool

OpenStudy (anonymous):

noooo lolll i didnt submit it

OpenStudy (anonymous):

right, now for the inside pool perimeter \[2(x^4+x^3-x)+2(2x^5+3x)\] then we have to subtract lets see what we get for the inside

OpenStudy (anonymous):

okay lets substract now ( :

OpenStudy (anonymous):

inside is \[4 x^5+2 x^4+2 x^3+4 x\]

OpenStudy (anonymous):

\[6 x^5+6 x^4-4 x^2+4 x-(4 x^5+2 x^4+2 x^3+4 x)\]for our final answer

OpenStudy (anonymous):

i mean we have to compute that one -hold on

OpenStudy (anonymous):

so is the final answer for question 2 ( :

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=6+x^5%2B6+x^4-4+x^2%2B4+x-%284+x^5%2B2+x^4%2B2+x^3%2B4+x%29 \[2 x^5+4 x^4-2 x^3-4 x^2\] is what i get

OpenStudy (anonymous):

alright same here but a plus sign guess i compute it wrong